OpenStudy (anonymous):

2(x+6) - (x+1) = 16 A. 3 B. 5 C. 2 D. 1

OpenStudy (anonymous):

you should remove the brackets by expanding the Left hand side. can you do this?

OpenStudy (anonymous):

ummm i think

OpenStudy (anonymous):

try it out

OpenStudy (anonymous):

2(x + 6) - (x + 1) = 16 2x + 12 - x - 1 = 16 2x - x = 16 - 12 + 1 x = 5

OpenStudy (anonymous):

thxs! i got that to :)

OpenStudy (anonymous):

well done